# Quasi-Dirac fermion: A source of neutrino mass and dark matter

### v1

@@ II Quasi-Dirac scheme
$m^{2}_{H^{-}}$ $=$ $\mu^{2}_{2}+\lambda_{3}v^{2}/2,$ (8)
$m^{2}_{S}$ $=$ $\mu^{2}_{2}+(\lambda_{3}+\lambda_{4}+\lambda_{5})v^{2}/2,$ (9)
$m^{2}_{A}$ $=$ $\mu^{2}_{2}+(\lambda_{3}+\lambda_{4}-\lambda_{5})v^{2}/2.$ (10)
@@
    Our neutrino mass generation scheme reveals a novel type of radiative inverse seesaw for which the radiative correction is directly related to neutrino mass, not to small Majorana masses $\mu,\kappa$ as of the normal scene [26, 27]. Additionally, while our scheme inherits a compelling feature as mass correction induced by scotogenic (i.e., dark) fields, it yields a radiative inverse-seesaw neutrino mass opposite to the scotogenic setup which generally gives a radiative canonical-seesaw neutrino mass, by contrast.
@@ III Dark matter and related phenomena
Figure 2: Annihilation processes of AA’s that set the dark matter abundance.
This model predicts the existence of a scalar dark matter, which is assumed to be $A$, i.e. $m_{A}<m_{S}$, $m_{A}<m_{H^{-}}$, and $m_{A}<m_{N_{1,2}}$, implying $\lambda_{5}>\mathrm{Max}(\lambda_{4},0)$ and $\bar{M}<M$. In the early universe, $A$ annihilates to usual particles via the gauge and Higgs portals, as depicted in Fig. 2, where there would exist a $u$-channel diagram to $ZZ$ associated with $t$-channel diagram to $ZZ$, which is skipped for brevity. Further, the $s$-channel diagram by $H$ portal may produce other particle pairs of the standard model, but their contributions are radically smaller that of $tt^{c}$ and are skipped too.
@@
$\mathcal{L}$ $\supset$ $-\frac{m_{t}}{v}\bar{t}tH-\frac{\lambda_{3}}{4}\left(2vH+H^{2}\right)A^{2}+\frac{g}{2c_{W}}Z^{\mu}S\stackrel{{\scriptstyle\leftrightarrow}}{{\partial}}_{\mu}A$ (15)
@@
$\langle\sigma v\rangle\simeq\frac{\lambda^{2}_{3}m^{2}_{t}}{64\pi m^{4}_{A}}\left(1-\frac{m^{2}_{t}}{m^{2}_{A}}\right)^{3/2}+\frac{g^{4}}{256\pi c^{4}_{W}m^{2}_{A}}+\frac{g^{4}}{128\pi m^{2}_{A}}+\frac{\lambda^{2}_{3}}{64\pi m^{2}_{A}},$ (16)
@@
$\langle\sigma v\rangle\simeq\left[\left(\frac{567.8\ \mathrm{GeV}}{m_{A}}\right)^{2}+\left(\frac{1.39\lambda_{3}\ \mathrm{TeV}}{m_{A}}\right)^{2}\right]\times 1\ \mathrm{pb}.$ (17)
Here note that the correct abundance $\Omega_{A}h^{2}\simeq 0.1\ \mathrm{pb}/\langle\sigma v\rangle\simeq 0.11$ requires $\langle\sigma v\rangle\simeq 1$ pb, and we use $\alpha=e^{2}/4\pi=1/128$, $g=e/s_{W}$, $s^{2}_{W}=0.231$. It is clear that (i) if $\lambda_{3}\ll g^{2}$, the gauge portal dominates the dark matter annihilation, and the correct abundance is set for $m_{A}=567.8$ GeV, (ii) if $\lambda_{3}\gg g^{2}$, the Higgs portal dominates the dark matter annihilation, and the correct abundance is set for $m_{A}=1.39\lambda_{3}$ TeV, and (iii) if $\lambda_{3}\sim g^{2}$, the correct relic density is set for $m_{A}\sim 816$ GeV, where both the portals give significant contributions.
@@
where $\lambda_{q}=\lambda_{3}m_{q}/2m_{A}m^{2}_{H}$. Hence, the dark matter ($A$) and nucleon ($\mathcal{N}=p,n$) scattering cross section is given by [29]
@@
where the reduced mass is $m_{r}=m_{A}m_{\mathcal{N}}/(m_{A}+m_{\mathcal{N}})\simeq m_{\mathcal{N}}\simeq 1$ GeV, while the nucleon coupling is summed over those at quark level interactions with respective nucleon form factors to be $\lambda_{\mathcal{N}}/m_{\mathcal{N}}\simeq 0.35\lambda_{3}/(2m_{A}m^{2}_{H})$ [30]. Taking $m_{H}=125$ GeV, we have
$\sigma^{\mathrm{SI}}_{A-\mathcal{N}}\simeq\left(\frac{25\lambda_{3}\ \mathrm{TeV}}{m_{A}}\right)^{2}\times 10^{-46}\ \mathrm{cm}^{2}.$ (20)
This prediction is appropriate to the experimental limit, i.e. $\sigma^{\mathrm{SI}}_{A-\mathcal{N}}=10^{-46}$$10^{-45}\ \mathrm{cm}^{2}$ according to $m_{A}=400$–1400 GeV, respectively [31], only for the case (i) above, which requires $\lambda_{3}\ll g^{2}$, thus matching $m_{A}=567.8\ \mathrm{GeV}=(25\lambda_{3}/\sqrt{1.3})\ \mathrm{TeV}$. Here note that $\sigma^{\mathrm{SI}}_{A-\mathcal{N}}\simeq 1.3\times 10^{-46}\ \mathrm{cm}^{2}$ for $m_{A}=567.8$ GeV [31]. It leads to $\lambda_{3}\simeq 0.025$, as expected. That said, the relic density is governed by the gauge portal, implying a dark matter mass $m_{A}=567.8$ GeV, while the direct detection cross section is set by the Higgs portal, requiring a dark matter vs. Higgs coupling $\lambda_{3}\simeq 0.025$.
@@ Acknowledgement
This research is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.01-2023.50.

### v2

@@ II Quasi-Dirac scheme
$m^{2}_{H^{-}}$ $=$ $\mu^{2}_{2}+(\lambda_{3}+\lambda_{4})v^{2}/2,$ (8)
$m^{2}_{S}$ $=$ $\mu^{2}_{2}+(\lambda_{3}+\lambda_{5})v^{2}/2,$ (9)
$m^{2}_{A}$ $=$ $\mu^{2}_{2}+(\lambda_{3}-\lambda_{5})v^{2}/2.$ (10)
@@
    Our neutrino mass generation scheme reveals a novel type of radiative inverse seesaw for which the radiative correction is directly related to neutrino mass, not to small Majorana masses $\mu,\kappa$ as of the normal scene [26, 27]. Additionally, while our scheme inherits a compelling feature as mass correction induced by scotogenic (i.e., dark) fields, it yields a radiative inverse-seesaw neutrino mass opposite to the scotogenic setup which generally gives a radiative canonical-seesaw neutrino mass, by contrast. That said, the dark fermions that couple to usual neutrinos may be dark matter, whereas those (i.e., vectorlike neutral fermions) in radiative inverse seesaw may decay to usual neutrinos and ordinary Higgs boson.
@@
Reference [29] has generalized the minimal scotogenic scheme by including an arbitrary number of inert scalar doublets, say $\eta$’s, and another arbitrary number of right-handed fermion singlets, say $N$’s, in which all $\eta$’s and $N$’s are odd under a $Z_{2}$ parity. The work has discussed nontrivial contributions to neutrino mass coming from the inert doublets $\eta$’s with a possibility that their small couplings with usual Higgs fields ($\lambda_{5}$’s) and the $Z_{2}$ symmetry are maintained at high energy. However, the right-handed fermion singlets have been assumed to be flavor diagonal, i.e. they have Majorana masses by themselves, for which the contributions of such right-handed fermion singlets to neutrino mass are simply added up from each individual contribution, as in the minimal scotogenic setup for three right-handed fermion singlets. At the first look, our proposal simply includes three $N_{R}$’s and three $N^{c}_{L}$’s, which are just six right-handed fermion singlets as proposed in [29]. However, the difference lies at global lepton number assignment, the original scotogenic model and its generalization in [29] put $L(N)=0$ for all $N$’s and $L(\eta)=-1$ for all $\eta$’s, while our model assigns $L(N)=1$ for all $N$’s and $L(\eta)=0$ for the inert scalar doublet, by contrast, which we recast it as a lepton-like charge, as supplied above, since this new charge might explain the $Z_{2}$ parity by gauge completion, discussed at the end. Interestingly, our proposal gives rise to novel physical results, because of the interplay of lepton-like symmetry and of pure-Majarona versus quasi-Dirac mass question. First of all, the lepton-like charge conservation demands that the Yukawa couplings of $N^{c}_{L}$’s to usual neutrinos $\nu_{L}$’s which violate this charge are small, i.e. $h^{\prime}\ll h$, as given above, which are different from those in [29] for which both $h,h^{\prime}$ conserve lepton number. Second, there is no symmetry reason that forbids $N$’s from obtaining a Dirac mass in both approaches, in our setup $N_{L,R}$ gain a Dirac mass as $M\bar{N}_{L}N_{R}$, while in previous setup [29], the Dirac mass is $M\bar{N}_{i}N^{c}_{j}$, where $N^{c}_{j}$ can be identified as the left-handed partner of the right-handed fermion $N_{i}$, for $i\neq j$. That said, in the setup [29], the $\lambda_{5}$ couplings among inert scalars and usual Higgs fields were small due to the violation of lepton number, but the pure-Majorana masses, say $m_{N_{i}}\bar{N}_{i}N^{c}_{i}$, conserved this lepton number to be large, which unfortunately did not exclude the Dirac mass $M\bar{N}_{i}N^{c}_{j}$, by contradiction. In other words, the hypothesis of pure-Majorana masses for $N$’s is unreasonable in any case. In our setup, the masses of $N$’s are determined by the lepton-like symmetry, namely the pure-Majorana masses $\mu_{L}\bar{N}_{L}N^{c}_{L}$ and $\mu_{R}\bar{N}^{c}_{R}N_{R}$ which violate the lepton-like symmetry must be small, compared to the Dirac mass $M\bar{N}_{L}N_{R}$, that is $\mu_{L,R}\ll M$, as the Yukawa couplings are, i.e. $h^{\prime}\ll h$, by the ’t Hooft’s naturalness [30]. Here, the Dirac mass $M$ conserves lepton-like (even lepton) charge, which would be large. This gives rise to quasi-Dirac fermions in our setup. Intriguingly, the quasi-Dirac fermions give nearly degenerate (and opposite) Majorana masses, i.e. $M$ and M-M, for which their contributions to neutrino mass are cancelled out. What results is a novel neutrino mass suppressed by $\mu_{L,R}/M$, as given above, which is a new observation of this work.
@@ III Dark matter and related phenomena
Figure 2: Annihilation processes of $A$’s that set the dark matter abundance.
This model predicts the existence of a scalar dark matter, which is assumed to be $A$, i.e. $m_{A}<m_{S}$, $m_{A}<m_{H^{-}}$, and $m_{A}<m_{N_{1,2}}$, implying $\lambda_{5}>\mathrm{Max}(-\lambda_{4},0)$ and $\bar{M}<M$. In the early universe, $A$ annihilates to usual particles via the gauge and Higgs portals, as depicted in Fig. 2, where there would exist a $u$-channel diagram to $ZZ$ associated with $t$-channel diagram to $ZZ$, which is skipped for brevity. Further, the $s$-channel diagram by $H$ portal may produce other particle pairs of the standard model, but their contributions are radically smaller that of $tt^{c}$ and are skipped too.
@@
$\mathcal{L}$ $\supset$ $-\frac{m_{t}}{v}\bar{t}tH-\frac{\lambda_{3}-\lambda_{5}}{4}\left(2vH+H^{2}\right)A^{2}+\frac{g}{2c_{W}}Z^{\mu}S\stackrel{{\scriptstyle\leftrightarrow}}{{\partial}}_{\mu}A$ (15)
@@
$\langle\sigma v\rangle\simeq\frac{(\lambda_{3}-\lambda_{5})^{2}m^{2}_{t}}{64\pi m^{4}_{A}}\left(1-\frac{m^{2}_{t}}{m^{2}_{A}}\right)^{3/2}+\frac{g^{4}}{256\pi c^{4}_{W}m^{2}_{A}}+\frac{g^{4}}{128\pi m^{2}_{A}}+\frac{(\lambda_{3}-\lambda_{5})^{2}}{64\pi m^{2}_{A}},$ (16)
@@
$\langle\sigma v\rangle\simeq\left[\left(\frac{567.8\ \mathrm{GeV}}{m_{A}}\right)^{2}+\left(\frac{1.39(\lambda_{3}-\lambda_{5})\ \mathrm{TeV}}{m_{A}}\right)^{2}\right]\times 1\ \mathrm{pb}.$ (17)
Here note that the correct abundance $\Omega_{A}h^{2}\simeq 0.1\ \mathrm{pb}/\langle\sigma v\rangle\simeq 0.11$ requires $\langle\sigma v\rangle\simeq 1$ pb, and we use $\alpha=e^{2}/4\pi=1/128$, $g=e/s_{W}$, $s^{2}_{W}=0.231$. It is clear that (i) if $\lambda_{3}-\lambda_{5}\ll g^{2}$, the gauge portal dominates the dark matter annihilation, and the correct abundance is set for $m_{A}=567.8$ GeV, (ii) if $\lambda_{3}-\lambda_{5}\gg g^{2}$, the Higgs portal dominates the dark matter annihilation, and the correct abundance is set for $m_{A}=1.39(\lambda_{3}-\lambda_{5})$ TeV, and (iii) if $\lambda_{3}-\lambda_{5}\sim g^{2}$, the correct relic density is set for $m_{A}\sim 816$ GeV, where both the portals give significant contributions.
@@
where $\lambda_{q}=(\lambda_{3}-\lambda_{5})m_{q}/2m_{A}m^{2}_{H}$. Hence, the dark matter ($A$) and nucleon ($\mathcal{N}=p,n$) scattering cross section is given by [31]
@@
where the reduced mass is $m_{r}=m_{A}m_{\mathcal{N}}/(m_{A}+m_{\mathcal{N}})\simeq m_{\mathcal{N}}\simeq 1$ GeV, while the nucleon coupling is summed over those at quark level interactions with respective nucleon form factors to be $\lambda_{\mathcal{N}}/m_{\mathcal{N}}\simeq 0.35(\lambda_{3}-\lambda_{5})/(2m_{A}m^{2}_{H})$ [32]. Taking $m_{H}=125$ GeV, we have
$\sigma^{\mathrm{SI}}_{A-\mathcal{N}}\simeq\left[\frac{25(\lambda_{3}-\lambda_{5})\ \mathrm{TeV}}{m_{A}}\right]^{2}\times 10^{-46}\ \mathrm{cm}^{2}.$ (20)
This prediction is appropriate to the experimental limit, i.e. $\sigma^{\mathrm{SI}}_{A-\mathcal{N}}=10^{-46}$$10^{-45}\ \mathrm{cm}^{2}$ according to $m_{A}=400$–1400 GeV, respectively [33], only for the case (i) above, which requires $\lambda_{3}-\lambda_{5}\ll g^{2}$, thus matching $m_{A}=567.8\ \mathrm{GeV}=[25(\lambda_{3}-\lambda_{5})/\sqrt{1.3}]\ \mathrm{TeV}$. Here note that $\sigma^{\mathrm{SI}}_{A-\mathcal{N}}\simeq 1.3\times 10^{-46}\ \mathrm{cm}^{2}$ for $m_{A}=567.8$ GeV [33]. It leads to $\lambda_{3}-\lambda_{5}\simeq 0.025$, as expected. That said, the relic density is governed by the gauge portal, implying a dark matter mass $m_{A}=567.8$ GeV, while the direct detection cross section is set by the Higgs portal, requiring a dark matter vs. Higgs coupling $\lambda_{3}-\lambda_{5}\simeq 0.025$. [^3] Of course, this value is appropriate to the magnitude of $\lambda_{5}\sim 0.01$ too.
@@ Acknowledgement
This research is funded by Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 103.01-2023.50. We thank Dr. Tran Van Que for a useful comment.